Circular No. 7765

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001hf IN MCG –03-23-17 M. Schwartz, Cottage Grove, OR; and W. D. Li, University of Califor-

nia at Berkeley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova (mag ∼ 16.1) on unfiltered images taken with the 0.5-m Tenagra III automated telescope on Nov. 27.3 and Dec. 4.5 UT. The new object is located at α = 9^{h}05^{m}07^{s}.44, δ = −18^{o}31′04′′.8 (equinox 2000.0), which is 4′′.5 west and 6′′.9 north of the nucleus of MCG –03-23-17. A Tenagra III image taken on Nov. 16.4 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2001gd IN NGC 5033 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2001gd (cf. IAUC 7761), obtained by P. Berlind on Dec. 4.52 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-IIb supernova well past maximum light: “The spectrum is almost identical to one of SN 1993J obtained on day 93 after explosion (see Matheson et al. 2000, A.J. 120, 1487). The helium lines are weak, but still discernable. The nebular-phase lines of magnesium, oxygen, and calcium are beginning to dominate the spectrum.”

SUPERNOVA 2001gc IN UGC 3375 M. Villi, Piacenza, Italy, reports that a CCD image taken with a 0.50-

m telescope on Nov. 17.30 UT in the course of the CROSS program (cf. IAUC 7373) shows SN 2001gc (IAUC 7759) at mag 17.1.

2001 QT_{297}

J. L. Elliot, Massachusetts Institute of Technology (MIT), reports that all three frames of 2001 QT_{297} recorded on Oct. 11 UT (all with the Sloan r′ filter) by D. J. Osip (MIT) with the MagIC camera on the Magellan I telescope (cf. IAUC 7733) have been analyzed further, yielding the following averages (for the secondary relative to the primary): difference in magnitude 0.70 ± 0.20; separation 0′′.61 ± 0′′.01; p.a. 116^{o}.8 ± 0^{o}.2 (measured north through east, as opposed to the non-standard north-through-west position angle reported on IAUC 7733).

2001 December 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green